OpenStudy (anonymous):

3. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 3x + 11y = 34.5 9x – 7y = -16.5

OpenStudy (anonymous):

\[3x+11y=34.5\]\[x=\frac{34.5-11y}{3}\] \[9x – 7y = -16.5\]\[9(\frac{34.5-11y}{3}) – 7y = -16.5\]\[103.5-33y – 7y = -16.5\]\[-40y = -120\]\[y = 3\]\[x = 0.5\]